# GAST: Sequential Gaussian Avatars with Hierarchical Spatio-temporal Context

### v1

@@ Abstract
3D human avatars, through the use of canonical radiance fields and per-frame observed warping, enable high-fidelity rendering and animating. However, existing methods, which rely on either spatial SMPL(-X) poses or temporal embeddings, respectively suffer from coarse rendering quality or limited animation flexibility. To address these challenges, we propose GAST, a framework that unifies 3D human modeling with 3DGS by hierarchically integrating both spatial and temporal information. Specifically, we design a sequential conditioning framework for the non-rigid warping of the human body, under whose guidance more accurate 3D Gaussians can be obtained in the observation space. Moreover, the explicit properties of Gaussians allow us to embed richer sequential information, encompassing both the coarse sequence of human poses and finer per-vertex motion details. These sequence conditions are further sampled across different temporal scales, in a coarse-to-fine manner, ensuring unbiased inputs for non-rigid warping. Experimental results demonstrate that our method combined with hierarchical spatio-temporal modeling surpasses concurrent baselines, delivering both high-quality rendering and flexible animating capabilities.
@@ 1 Introduction
Figure 1: Illustration of the Hierarchical Spatio-temporal Context Design. We model the motion-dependent appearance variations with both coarse skeleton condition $\Delta P$ and fine-grained point-wise velocity condition $\mathcal{V}$. $\Delta P$ describes the overall human pose motion, which is derived from the difference between the human poses at adjacent frames. $\mathcal{V}$ is the point-wise velocity that indicates finer-grained motion in local regions. Based on $\Delta P$ and $\mathcal{V}$, we utilize an MLP to predict more accurate non-rigid deformation of each Gaussian primitive. To achieve robust deformation, we sample the overall motion trend and inter-frame details via diverse time intervals $s$ for $\Delta P$ and $\mathcal{V}$. The rendering results demonstrate that our method excels in regions with complex texture details as well as in areas where the appearance varies dramatically with motions.
@@
Recent research on digital humans has highlighted the efficiency of 3D Gaussian Splatting [23] (3DGS), demonstrating its capability for high-quality and real-time rendering. By defining T-pose human Gaussians in a canonical space and employing specific warping techniques, we can render human avatars under various observations. The canonical Gaussian units and warping weights are jointly optimized under supervision from human video input.
@@
The primary bottleneck in this task stems from the limited modeling of non-rigid warping, requiring a trade-off between rendering quality and animating capability. One stream of approaches [17, 44] introduces the prior from SMPL(-X) [37, 42] poses, leveraging human poses to guide motions of each observation and relying on Linear Blend Skinning (LBS) for motion warping. Although they are well-suited for pose-driven animation, they often struggle with per-frame non-rigid warping in scenarios with complex garments. Other approaches [33, 51] utilize temporal embeddings instead of SMPL poses to model each frame independently, achieving high-quality rendering but making it challenging to animate avatars.
We experimentally find that the existing human pose condition [17, 44] and temporal embeddings are insufficient to fully capture the expression of human motion. On the one hand, while the human pose represents the spatial information of a template body, it fails to capture local details (e.g., garment deformation). On the other hand, existing temporal embeddings [33, 51] maintain temporal continuity in rendering by sacrificing geometric awareness of the human body, limiting their effectiveness for animation. In this paper, we aim to strike an ideal balance in both rendering and animating of human avatars by introducing a hierarchical spatio-temporal context. To this end, we refine the input conditions for non-rigid warping by modeling a motion sequence, ensuring correspondence of Gaussians across different observations.
The desired motion sequence should encapsulate rich spatio-temporal information to effectively guide the non-rigid warping process. Previous NeRF[38]-based methods [5] construct the motion sequence with human pose residuals to represent human motion. However, the human pose, as an inherently global representation, struggles to capture finer motion details. The explicit structure of 3DGS enables localized temporal modeling, facilitating the construction of a hierarchical motion sequence by encoding motion residuals from template vertices in SMPL(-X). This coarse-to-fine, pose-to-vertex design (as illustrated in Fig. 1) for the motion sequence captures hierarchical spatial motion while guaranteeing temporal continuity. How to determine the optimal stride for motion sequence remains a significant challenge. The motion sequence with a larger stride captures longer motion but loses details between adjacent frames, whereas the motion sequence with a smaller stride covers a shorter motion range but retains richer inter-frame details. To further incorporate detailed human motion, we propose a multi-scale sampling strategy for the motion sequence. The final input to the non-rigid network consists of motion sequences sampled at multiple temporal scales, arranged in a coarse-to-fine manner. This hierarchical temporal design combines both motion range and inter-frame details, further enhancing the model’s generalization ability in rendering from novel views.
We test our pipeline, GAST, on I3D-Human dataset [5] and DNA-Rendering dataset [6], which contain complex garments motions beyond the scope of the template body. Our method outperforms other baselines on these datasets, achieving compatibility in both rendering and animating. To summarize, our contributions are three-fold as follows:
@@
    GAST, the first framework for sequential Gaussian avatars, achieving a balance between high-fidelity rendering and animation through spatio-temporal context;
@@
    a hierarchical motion sequence that combines global human pose and localized vertex residuals, guiding non-rigid warping from coarse to fine;
@@
    a multi-scale sampling strategy for selecting strides for motion sequences, integrating both broader motion dynamics and finer inter-frame details.
@@ 2 Related Work
2.1 Human Avatars Based on SMPL(-X)
Recent advances in digital human modeling have benefited greatly from the SMPL(-X) series [37, 42], which represents the human body by decomposing it into distinct shape components and pose-related motions, utilizing 3D mesh scanning techniques and principal component analysis (PCA). The pose blend shapes in SMPL(-X) enable human body deformation through a linear blend of poses of different joints, significantly improving the efficiency of animating digital humans. The widespread adoption of SMPL(-X) has further been boosted for human body animation, driven by methods [7, 46] that estimate parameters of SMPL(-X) from 2D image inputs. Despite its wide applications, the SMPL(-X) framework has its limitations. The skin-tight 3D meshes, based on pose blend shapes, struggle to model bodies with garments or accessories. This brings a challenge for complex human rendering based on SMPL(-X), particularly when dealing with garments and accessories.
2.2 SMPL(-X) Poses Guided Neural Human
Implicit neural methods, such as NeRF [38] and 3DGS [23], have become key contributors to digital human reconstruction due to their high-quality rendering capabilities. Some methods get inspiration from SMPL(-X) to build neural animatable 3D human avatars. These approaches typically map the T-pose to a canonical space and apply LBS weights to perform rigid transformations. Animation is then realized by mapping points from the canonical space to the observation space, with each frame’s human pose determining the transformation. NeRF-based approaches [49, 43, 25, 5, 14, 2, 4, 3, 9, 12, 13] focus on reconstructing human avatars from monocular or multi-view synchronized videos. For more accurate human geometry, [47] incorporate smooth priors from neural Signed Distance Functions (SDF). More recent advances [32, 35, 55, 28, 17, 44, 27, 18, 40, 22, 16, 29, 34, 53, 24, 39, 21, 54] leverage 3DGS, resulting in faster and high-fidelity human rendering. We also notice that some works [5, 18] model a human pose sequence as the context to improve rendering quality. But they suffer from low-speed rendering and lack fine details of motion caused by simply integrating coarse human pose embedding with NeRF.
2.3 Human Rendering with Temporal Embeddings
@@
The earliest modeling of dynamic radiance field [11, 30, 31, 50, 8, 1, 36, 41, 20, 10, 45] can be traced back to temporal embeddings in general scenes, where each frame’s observation is obtained by constructing a canonical space and a time-conditioned deformation field. Motivated by these, a stream of human researches [33, 51] focus on pure rendering quality, employing temporal embeddings instead of human pose to encode each frame of human videos. These methods often achieve high-quality, temporally continuous rendering. However, they struggle with the lack of geometric constraints from human pose, making parameterized animating unfeasible.
@@ 3 Preliminary
where $K$ is the total bone number and $\mathbf{B}_{k}$ is the transformation matrix for each bone. Specifically, $\mathbf{B}_{k}$ consists of a rotation and translation matrix $\mathbf{R}_{k}$ and $\mathbf{b}_{k}$. Here, $\mathbf{R}_{k}$ represents the global rotation of each joint, influenced by the local joint rotations $\bm{r}$, while $\mathbf{b}_{k}$ denotes each joint’s translation, determined by the joint positions $\bm{j}$ and the human shape parameters $\beta$. Note that the linear blending weight $\omega^{k}$ depends on $\mathbf{x}$ and is regressed from comprehensive human meshes.
3D Gaussian Splatting [23] utilizes a set of point-based Gaussian primitives to explicitly represent scenes, enabling real-time and high-quality rendering. Each Gaussian primitive is defined by its center position $\mathbf{x}\in\mathbb{R}^{3}$ and covariance matrix $\mathbf{\Sigma}\in\mathbb{R}^{3\times 3}$.
To ensure positive semi-definiteness and simplify optimization, the covariance matrix $\Sigma$ is decomposed into a rotation matrix $\mathbf{R}$ and a scaling matrix $\mathbf{S}$ following
@@
After projection, the pixel color $C$ is computed by blending $N$ ordered Gaussian primitives overlapping at the pixel as
@@
$L_{photo}=\lambda_{1}L_{1}(\hat{I},I)+(1-\lambda_{1})(1-SSIM(\hat{I},I)),$ (5)
where $\hat{I}$ and $I$ denote the rendered and ground-truth images, and $\lambda_{1}$ controls the balance between the two terms.
@@ 4 Method
Fig. 2 shows the framework of our method. We use the explicit point-based 3DGS as the representation of the human body. Given a collection of input cameras and images, we optimize a set of Gaussian primitives $\{\mathcal{G}_{i}\}_{i=1}^{i=n}$ to fit the body’s shape and appearance. Each Gaussian primitive $\mathcal{G}_{i}$ includes the center position $\mathbf{x}$, scaling vector $\mathbf{s}$, rotation quaternion vector $\mathbf{r}$, color feature $sh$ and opacity $\mathbf{\alpha}$, where $\mathbf{x}$ is initialized from the SMPL template vertices $\{\mathbf{T}_{i}\}_{i=1}^{i=N}$. During the optimization process, we first apply non-rigid deformation with hierarchical spatio-temporal conditions to capture more accurate and fine-grained details (Sec. 4.1). Next, we adopt Linear Blend Skinning (LBS) to map canonical Gaussian primitives $\mathcal{G}$ to observation space and render images via differentiable splitting (Sec.4.3).
Figure 2: Overview of GAST. We first initialize canonical Gaussian positions $\mathbf{x}$ with SMPL template vertexes. We propose a hierarchical spatio-temporal context for non-rigid deformation to better capture complex human motion. The skeleton motion condition $f_{\Delta\mathcal{P}}$ aims to distinguish the appearance variation caused by overall body movement. The point motion condition $f_{\mathcal{V}}$, sampled from the vertex motion template $\mathcal{F}_{\mathcal{V}}$ (points in different colors represent different motions), is designed to capture finer-grained temporal information in local regions. We utilize an MLP $\mathcal{E}_{non-rigid}$ to predict the non-rigid deformation of each Gaussian based on both $f_{\Delta\mathcal{P}}$ and $f_{\mathcal{V}}$ conditions. The non-rigid deformed Gaussians are then warped into observation space via the standard LBS transformation for rendering.
4.1 Non-Rigid Deformation with Hierarchical Spatio-temporal Context
Coarse Skeleton Motion. As illustrated in Fig. 1, we introduce a coarse-to-fine spatio-temporal context to excavate additional conditions for non-rigid deformation. Typically, most human 3D avatar methods use the current frame pose as a condition to predict non-rigid deformation [17, 44]. While this provides spatial information to distinguish the motion of different body parts, it fails to capture temporal motion changes, as discussed in [5]. Therefore, given a frame at time $t$, we consider a sequence of regularly interval-sampled frames $\mathcal{T}$ to model inter-frame body motion variations, defined as
@@ 4.1 Non-Rigid Deformation with Hierarchical Spatio-temporal Context
where $L$ is the sequence length and $s$ is the time interval. The coarse motion of body skeletons at each time step $t\in\mathcal{T}$ can be derived by calculating the difference of poses between adjacent frames (as illustrated in Fig. 1), following
@@
where $P\in\mathbb{R}^{K\times 3}$ is the body pose and $\delta$ denotes the difference $\Delta\mathcal{P}$ between two poses in axis-angle form. We employ an MLP $\mathcal{E}_{\Delta\mathcal{P}}$ to encode the sequential skeleton motion $\Delta\mathcal{P}\in\mathbb{R}^{L\times K\times 3}$, flattened along the temporal dimension, into a skeleton motion embedding $f_{\Delta\mathcal{P}}\in\mathbb{R}^{32}$, which serves as a condition for the subsequent non-rigid deformation. Further details are provided in the supplementary material.
@@
We sample the velocity from this motion template field $\mathcal{F}_{\mathbf{V}}$ for each Gaussian primitive. To obtain robust sampled velocities and model body’s local region information, we input the $\tau$ nearest template vertexes’ velocities into an MLP $\mathcal{E}_{knn}$ to learn a local velocity embedding

### v2

# Sequential Gaussian Avatars with Hierarchical Motion Context

@@ Abstract
The emergence of neural rendering has significantly advanced the rendering quality of 3D human avatars, with the recently popular 3DGS technique enabling real-time performance. However, SMPL-driven 3DGS human avatars still struggle to capture fine appearance details due to the complex mapping from pose to appearance during fitting. In this paper, we propose SeqAvatar, which excavates the explicit 3DGS representation to better model human avatars based on a hierarchical motion context. Specifically, we utilize a coarse-to-fine motion conditions that incorporate both the overall human skeleton and fine-grained vertex motions for non-rigid deformation. To enhance the robustness of the proposed motion conditions, we adopt a spatio-temporal multi-scale sampling strategy to hierarchically integrate more motion clues to model human avatars. Extensive experiments demonstrate that our method significantly outperforms 3DGS-based approaches and renders human avatars orders of magnitude faster than the latest NeRF-based models that incorporate temporal context, all while delivering performance that is at least comparable or even superior. Project page: https://zezeaaa.github.io/projects/SeqAvatar/
Figure 1: Illustration of the Hierarchical Motion Context Design. We model the motion-dependent appearance variations with both coarse skeleton condition $\Delta P$ and fine-grained point-wise velocity condition $\mathcal{V}$. $\Delta P$ describes the overall human skeleton motion, which is derived from the difference between the human poses at adjacent frames. $\mathcal{V}$ is the point-wise velocity that indicates finer-grained motion in local regions. To achieve robust deformation, we propose Spatio-temporal Multi-scale Sampling, which samples the overall motion trend and inter-frame details via diverse time intervals $s$ for $\Delta P$ and $\mathcal{V}$.
@@ 1 Introduction
Recent research on digital humans has highlighted the efficiency of 3D Gaussian Splatting (3DGS) [27], demonstrating its capability for high-quality and real-time rendering. By defining T-pose human Gaussians in a canonical space and employing specific warping techniques, we can render human avatars from diverse perspectives and in any pose. The canonical Gaussian primitives and warping weights are then jointly optimized under supervision from video inputs.
@@
Animatable human avatar reconstruction is primarily hindered by limited modeling of non-rigid warping, forcing a trade-off between rendering quality and animation capability. Current methods  [22, 56] incorporate SMPL(-X) pose priors  [43, 51] to guide motion for each observation and rely on Linear Blend Skinning (LBS) for warping. Although these approaches are effective for pose-driven animation, they often struggle with per-frame non-rigid warping in scenarios involving complex garments.
We experimentally find that current human pose conditions [22, 56] do not fully capture the complex one-to-many mapping from pose to appearance, particularly for 3DGS-based approaches [27]. These methods [68, 56] rely on the spatial information of a template body in each frame’s human pose to predict non-rigid deformations. However, they often overlook local details, such as garment deformations far from the skeleton, and cannot resolve cases where the same pose corresponds to different appearances during complex motions. Furthermore, although previous NeRF-based [44] method [8] has attempted to model motion sequences using human pose residuals, the inherently global nature of the pose sequence limits the ability to capture finer motion details. Naive pose sequence modeling does not fully account for the explicit characteristics of 3DGS.
In this paper, we introduce a hierarchical motion context condition for 3DGS-based human avatar modeling to address the limitations of relying solely on human pose, which provides only limited global skeletal information. Our approach improves the ability of 3DGS-based methods to accurately capture the complex relationship between human pose and appearance in challenging scenarios. Specifically, we design a coarse-to-fine motion condition that incorporates both overall skeletal movements and fine-grained point-wise motions. Leveraging the explicit nature of Gaussian primitives, this condition seamlessly integrates into 3DGS-based methods, enabling more precise predictions of complex non-rigid deformations. To further enhance robustness, we propose a spatio-temporal multi-scale sampling strategy for constructing the motion context with a larger receptive field. Spatially, we consider the motion states of neighboring points within the local region of each Gaussian primitive to obtain more stable motion embeddings. Temporally, we capture human motion patterns across multiple time scales, combining long-term trends with fine-grained inter-frame motion details. This improves the model’s generalization to complex human movements.
To summarize, our key contributions are as follows:
@@
    We propose a novel hierarchical motion condition that integrates coarse-to-fine human motion, combining global skeletal poses with localized vertex residuals to enhance non-rigid deformation prediction.
@@
    We introduce a spatiotemporal multiscale sampling strategy that expands the receptive field of hierarchical motion context, improving generalization to complex motions.
@@
    Experiments on the I3D-Human [8], DNA-Rendering [9], and ZJU-MoCap [53] datasets show the effectiveness of the proposed SeqAvatar, which is capable of modeling details of the human body in complex motions.
@@ 2 Related Work
Neural Rendering. Neural rendering techniques have brought significant progress to human reconstruction and rendering. In particular, Neural Radiance Fields (NeRF) [44] introduces an implicit scene representation that models color and density using multilayer perceptrons, delivering photorealistic rendering results. More recently, point-based 3D Gaussian Splatting [27] utilizes 3D Gaussians to represent scenes explicitly, achieving real-time high-quality rendering. Building upon NeRF and 3DGS, subsequent works have advanced neural rendering across various dimensions, e.g., enhancing visual fidelity [1, 2], improving sparse-view reconstruction quality [10, 63, 80, 71, 52], enabling pose-free optimization [67, 37, 61, 58], modeling dynamic scenes [55, 13, 34, 49, 69, 74], and accelerating training and inference [18, 46, 12, 60, 4, 39]. These developments also benefit human avatar modeling, which demands high-quality and efficient rendering.
SMPL(-X)-Based Neural Human Modeling. The SMPL(-X) family [43, 51] provides a parametric representation of the human body by decomposing it into pose-related and shape-related components using 3D mesh scans and principal component analysis (PCA). Its pose blend shapes enables body deformation through joint-wise pose blending, offering an efficient and compact way for animation. SMPL(-X) has thus become a cornerstone in human body modeling and animation, with many methods [11, 59] estimating its parameters directly from 2D inputs. Recent neural human reconstruction methods integrate SMPL(-X) with implicit representations such as NeRF [44] and 3DGS [27] to enable animatable avatars with high rendering quality. Different from purely 2D image animation methods [73, 65, 62, 48], these methods typically register the input data to a canonical T-pose space and use linear blend skinning (LBS) to transform 3D points into observation space based on SMPL(-X) poses. NeRF-based approaches [68, 53, 30, 8, 20, 5, 7, 6, 14, 17, 19, 64, 75] focus on monocular or multi-view reconstruction. More recently, 3DGS-based methods [36, 41, 81, 32, 22, 56, 31, 23, 47, 26, 21, 33, 40, 78, 29, 45, 25, 79, 76] have gained popularity due to their real-time rendering and high fidelity. Some works [8, 23] further incorporate pose sequences as temporal context, but they suffer from slow rendering and lack fine details of motion caused by simply integrating coarse human pose embedding with NeRF.
@@
Human Rendering with Temporal Embeddings The earliest modeling of dynamic radiance field [16, 34, 35, 70, 13, 3, 42, 50, 24, 15, 57] can be traced back to temporal embeddings in general scenes, where each frame’s observation is obtained by constructing a canonical space and a time-conditioned deformation field. Motivated by these, a stream of research [38, 72] focuses on pure rendering quality, employing temporal embeddings instead of human pose to encode each frame of human videos. Although these methods achieve high-quality, temporally continuous rendering, they struggle with the lack of geometric constraints from human pose, making parameterized animation unfeasible.
@@ 3 Preliminary
where $K$ is the total bone number and $\mathbf{B}_{k}$ is the transformation matrix for each bone. Specifically, $\mathbf{B}_{k}$ consists of a rotation and translation matrix $\mathbf{R}_{k}$ and $\mathbf{b}_{k}$. Here, $\mathbf{R}_{k}$ represents the global rotation of each joint, influenced by the local joint rotations $\bm{r}$, while $\mathbf{b}_{k}$ denotes each joint’s translation, determined by the joint positions $\bm{j}$ and the human shape parameters $\beta$. The linear blending weight $\omega^{k}$ depends on $\mathbf{x}$ and is regressed from comprehensive human meshes.
3D Gaussian Splatting [27] utilizes a set of point-based Gaussian primitives to explicitly represent scenes, enabling real-time and high-quality rendering. Each Gaussian primitive is defined by its center position $\mathbf{x}\in\mathbb{R}^{3}$ and covariance matrix $\mathbf{\Sigma}\in\mathbb{R}^{3\times 3}$. To ensure positive semi-definiteness and simplify optimization, the covariance matrix $\Sigma$ is decomposed into a rotation matrix $\mathbf{R}$ and a scaling matrix $\mathbf{S}$:
@@
After projection, the pixel color $C$ is computed by blending $N$ ordered Gaussian primitives overlapping at the pixel:
@@
$L_{photo}=\lambda L_{1}(\hat{I},I)+(1-\lambda)(1-SSIM(\hat{I},I)),$ (5)
where $\hat{I}$ and $I$ denote the rendered and ground-truth images, and $\lambda$ controls the balance between the two terms.
@@ 4 Method
Fig. 2 shows the framework of our method. We use the explicit point-based 3DGS as the representation of the human body. Given a collection of input cameras and images, we optimize a set of Gaussian primitives $\{\mathcal{G}_{i}\}_{i=1}^{i=n}$ to fit the body’s shape and appearance. Each Gaussian primitive $\mathcal{G}_{i}$ includes the center position $\mathbf{x}$, scaling vector $\mathbf{s}$, rotation quaternion vector $\mathbf{r}$, color feature $sh$ and opacity $\mathbf{\alpha}$, where $\mathbf{x}$ is initialized from the SMPL template vertices $\{\mathbf{T}_{i}\}_{i=1}^{i=N}$. During the optimization process, we first apply a coarse-to-fine motion context to capture more accurate and fine-grained details (Sec. 4.1). To mitigate overfitting, we propose Spatio-Temporal Multi-Scale Sampling to obtain more robust point-wise Gaussian motion, which serves as the embedding for non-rigid deformation (Sec.4.2). Next, we adopt Linear Blend Skinning (LBS) to map canonical Gaussian primitives $\mathcal{G}$ to observation space and render images via differentiable splitting (Sec.4.3).
Figure 2: Overview of the proposed method. We first initialize canonical Gaussian positions $\mathbf{x}$ with SMPL template vertexes. For each Gaussian, we derive both coarse skeleton motion condition $f_{\Delta\mathcal{P}}$ and fine-grained vertex motion condition $f_{\mathcal{V}}$ that sampled from the vertex motion template $\mathcal{F}_{\mathcal{V}}$ (points in different colors represent different motions). Based on such hierarchical motion information, we utilize an MLP $\mathcal{E}_{non-rigid}$ to better predict each Gaussian’s non-rigid deformation prediction. The non-rigid deformed Gaussians are then warped into observation space via the standard LBS transformation for rendering.
4.1 Non-Rigid Deformation with Coarse-to-fine Motion Context
Coarse Skeleton Motion. As illustrated in Fig. 1, we introduce a coarse-to-fine motion context to excavate additional conditions for non-rigid deformation. Typically, most human 3D avatar methods use the current frame pose as a condition to predict non-rigid deformation [22, 56]. While this provides spatial information to distinguish the motion of different body parts, it fails to capture temporal motion changes, as discussed in [8]. Therefore, given a frame at time $t$, we consider a sequence of regularly interval-sampled frames $\mathcal{T}$ to model inter-frame body motion variations:
@@ 4.1 Non-Rigid Deformation with Coarse-to-fine Motion Context
where $L$ is the sequence length and $s$ is the time interval. The coarse motion of body skeletons at each time step $t\in\mathcal{T}$ can be derived by calculating the difference of poses between adjacent frames as illustrated in Fig. 1, following
@@
where $P\in\mathbb{R}^{K\times 3}$ is the body pose and $\delta$ denotes the difference $\Delta\mathcal{P}$ between two poses in axis-angle form. We employ an MLP $\mathcal{E}_{\Delta\mathcal{P}}$ to encode the sequential skeleton motion $\Delta\mathcal{P}\in\mathbb{R}^{L\times K\times 3}$, flattened along the temporal dimension, into a skeleton motion embedding $f_{\Delta\mathcal{P}}\in\mathbb{R}^{32}$,
$f_{\Delta\mathcal{P}}=\mathcal{E}_{\Delta\mathcal{P}}(\Delta\mathcal{P}),$ (8)
which serves as a condition for the subsequent non-rigid deformation. Further details are provided in the supplementary material.
@@
We then sample each Gaussian’s velocity from this motion template field $\mathcal{F}_{\mathbf{V}}$, which is discussed in Sec. 4.2.
4.2 Spatio-temporal Multi-scale Sampling (STMS)
To enhance the robustness of human motion conditions, we propose modeling both local region motion information of the human body and motion patterns across different temporal windows to mitigate overfitting and improve generalization by capturing more comprehensive motion dynamics.
In the spatial dimension, we sample the $\tau$ nearest template to model the body’s local region motion more robustly. Specifically, for each canonical Gaussian primitive $\mathcal{G}_{i}$, the $\tau$ nearest vertexes’ velocities are sampled as input to an MLP $\mathcal{E}_{knn}$ to learn a local point-wise motion embedding